OpenStudy (anonymous):

A right triangle has legs of length 7 and 9. What is the length of the hypotenuse?

OpenStudy (eric_d):

Welcome to Open Study !!! @BradSuga

OpenStudy (eric_d):

Use pythagorean theorem @BradSuga

OpenStudy (eric_d):

c^2= a^2+ b^2

OpenStudy (here_to_help15):

Ok so a^2+B^2=C^2

OpenStudy (here_to_help15):

7^2+9^2=c^2

OpenStudy (eric_d):



OpenStudy (here_to_help15):

so 7*7=49 and 9*9=81 So 81 +49 ______________ 130 then you square that \[\sqrt{130}\]

OpenStudy (anonymous):

What about the c^2 what exactly happens to the power?

OpenStudy (anonymous):

Ah but yet again you put it in radical

OpenStudy (anonymous):

The diameter of a circular cylindrical holding tank can be modeled by the function, d=2\/a/π , where A is the surface area of the circular base of the tank. If the surface area of the base is 64π square feet, what is the diameter of the tank?